OpenStudy (anonymous):

solve by completing the square 4x^2-4x-63=0

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

just want you to check my answer

OpenStudy (anonymous):

\[4x^2-4x-63=0\] \[4x^2-4x=63\] \[x^2-x=\frac{63}{4}\] \[(x-\frac{1}{2})^2=\frac{63}{4}+\frac{1}{4}=\frac{64}{4}=16\] \[x-\frac{1}{2}=\pm\sqrt{16}=\pm4\] \[x-\frac{1}{2}=4\iff x=\frac{9}{2}\] \[x-\frac{1}{2}=-4\iff x=-\frac{7}{2}\]

OpenStudy (anonymous):

\[(2x-9)(2x+7)=0\] may have been easier

OpenStudy (anonymous):

yeah but i was told to solve it by completing the square

OpenStudy (anonymous):

wait where did that 1/4 come from

OpenStudy (anonymous):

that you added to 63

OpenStudy (anonymous):

plz tell

OpenStudy (anonymous):

nvm just realized